OpenStudy (anonymous):

2. A survey asked 100 students about what meals they ate in the university cafeteria. The data were as follows:  48 ate breakfast  49 ate lunch  38 ate dinner  17 ate breakfast and lunch  15 ate breakfast and dinner  18 ate lunch and dinner  7 ate all three meals (a) How many students ate only breakfast in the cafeteria? (b) How many students ate at least two meals in the cafeteria? (c) How many students ate breakfast and lunch, but not dinner, in the cafeteria? (d) How many students did not eat any meals in the cafeteria?

OpenStudy (anonymous):

i had already answered this one josh

OpenStudy (anonymous):

i cant get c

OpenStudy (anonymous):

48-17-8= 23 49-17-8 = 24 so 47 for c ?

OpenStudy (anonymous):

for c 23+10+21 = 54. for d 100 - (23+10+21+8+7+11+12) = 100 - 92 = 8

OpenStudy (anonymous):

thanks so much, did you get my other problem i aasked ?

OpenStudy (anonymous):

no type it properly

OpenStudy (anonymous):

Using the simplex method, solve the following linear programming problem. minimize z = x1 + 2x2 - 5x3 subject to x1 + 4x2 + x3 >_ 25 x1 + 7x2 + 6x3 <_100 x1 >_0; x2 >_ 0; x3 >_ 0

OpenStudy (anonymous):

hey i will post the answer later on i have got some work will catch u in half an hour or so

OpenStudy (anonymous):

see here u need to apply Big M Method for solving this and u r required to introduce artificial variables

OpenStudy (anonymous):

its hard and i have class in 30 hour so no worries i have no clue

OpenStudy (anonymous):

this one was much better c u later

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hey see in this case u need to make pivot table wich i will be unable in making it here so will send u it in attachment